The Fermi enrgy of copper is 7 eV. (use h = 6.6 ×10−34JHz−1 , 1eV = 1.602×10−19J, me=9.1×10−31 kg ).
The Fermi momentum of electron in copper is?
1
Expert's answer
2021-09-21T11:50:33-0400

The electron has a kinetic energy equal to the Fermi energy for copper


"K=7\ (eV)=11.214\cdot10^{-19}\ (J)" . So,


"p=\sqrt{2mK}=\sqrt{2\cdot9.1\cdot10^{-31}\cdot11.214\cdot10^{-19}}=1.43\cdot10^{-24}\ (kg\cdot m/s)="


"=2.678\cdot10^3\ (eV/c)" .